1/x+1/(x+1)= 1/1.5
why can't i get this?

Answer by edjones(8007)

1/x+1/(x+1)= 1/1.5
1/x + 1/(x+1) = 2/3
We have to get rid of the fractions. Multiply both side times 3x(x+1)
3(x+1)+3x=2x(x+1) Much better!
3x+3+3x=2x^2+2x
6x+3=2x^2+2x
-2x^2+4x+3=0 subtract the right side from the left.
2x^2-4x-3=0 multiply both sides by -1.
x^2-2x-3/2=0 divide each side by 2.
x^2-2x =3/2
x^2-2x+1=3/2+1 complete the square
(x-1)^2=2.5
x-1=+-sqrt(2.5) sqrt of both sides
x=1+-sqrt(2.5)
x= 2.58114
x= -0.58114
Ed